OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2-6x+8=y A. (–3,– 6) B. (3,–1) C. (–3,1) D. (–1, 3)

OpenStudy (anonymous):

vertex is at \(x=-\frac{b}{2a}=-\frac{-6}{2\times 1}=3\)

OpenStudy (anonymous):

leaves only one choice

OpenStudy (anonymous):

so the answer is D?